Find remainder when the sum 4^37 + 6^37 is divided by 25.
A) 5 B) 10 C) 15 D) 20 E) None

Answer by edjones(8007)

4^37 + 6^37 = 10^37
10^37= 1 followed by 37 zeros
There is no remainder when divided by 25. E
.
Ed